

********************************************************************
UARS WEEKLY STATUS REPORT - 12 JUNE 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE persists with its customary operating status, continuing to amass
celestial statistics during all obtainable solar observation intervals and
astral calibration analyses, whenever feasible.
***********
SUSIM
***********
SUSIM has successfully gathered its normal solar, offset, lamp, and
occultation scan data without apparent problems.
Work is nearly finished on version 7 of the SUSIM production software.
***********
PEM
***********
PEM was not active.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
HALOE continues successful operations during this operational period.
HALOE will be powered off at end-of-day today for the upcoming yaw
maneuver. All engineering parameters are nominal and the science data
results are excellent.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI is currently off. Observations from May 31 to June 10 were normal.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
An orbit-adjust maneuver for the UARS spacecraft is being planned
for 12 August 1998. In preparation for the thruster firing on
12 August, small thruster "burns" for calibration are being
planned for 28 July 1998.
***********
CDHF
***********
The CDHF began receiving orbit data on a daily basis on 06/03/98.
This is a change from the old schedule of receiving this data
on Mon - Wed - Fri.
The CDHF has received a total of 393 CD-ROM requests as of 12-June-1998.
There were three orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
26 days PEM (3.32) PROD (1379R-1339C=40L)
81 days HAL (4.80) PROD (L1) (1868R-392C-2F=1474L)
30 days HAL (4.80) PROD (L3) (1868R-105C=1763L)
309 days SOL (4.61) PROD (L1) (2404R-1463C=941L)
76 days SOL (4.61) PROD (L2-L3)(2304R-245C=2059L)
The CDHF will load April 1998 current day DAAC data in June 1998.
CDHF pending software upgrades: IDL 5.1, OpenIngre
***********
FOT
***********
REPORTING PERIOD
This report covers 30 May 1998 (Orbit 36693, GMT Day 150) through 05 June
1998 (Orbit 36797, GMT Day 156).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 151/0135 36709 On
MLS 152/2330 36737 Off
HALOE 153/0107 36738 On
A summary of Instrument operations supplied by the Mission Planning Group
for the month of May 1998 is contained in Appendix F, UARS Instrument
Operations Summary.
The beta angle is now decreasing toward a minimum angle of 0.0 degrees on
15 June 1998 (DOY 166).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 25.4 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 30 May 1998 and 05 June 1998 are provided in Appendix D.
The clock error ranged from -10.9 to +9.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 02
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 01
% HRDI 02
% MLS 01
% HALOE 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 03
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/155 show the observatory data loss to be
20 hours, 43 minutes, 36.628 seconds (an increase of 6.976 seconds since
last report period). This is a 0.0353 percent data loss which equals a
99.9647 percent data capture for the mission.
The increase of 6.976 seconds occurred over a five day period:
7 0.512 seconds on Day 151 due to a hit in data,
7 0.384 seconds on Day 152 due to hits in data,
7 3.072 seconds on Day 153 due to hits in data,
7 2.016 seconds on Day 154 due to hits in data and line errors, and
7 0.992 seconds on Day 155 due to insufficient overlap.
The FOT made a few more modifications to the OBCSTF resident POCC database
this week. The latest RtestS version of Database 67 was generated this
week. The resulting database tape was given to the MSOCC software
coordinator, who subsequently installed the database onto the PE41 disk
pack. Database testing was performed using PE41, and no problems with the
database have been discovered to date.
The FOT built three HALOE RTS loads this week as requested by HALOE
operations to accommodate the HALOE instrument turn-on procedure. The
three RTS loads were thought (by HALOE operations) to already be available
for use, but as it turned out, the loads were sufficiently old and were
deleted from the CMS system via normal file maintenance.
New versions of RTSs 29 and 30 were generated this week. The purpose of
the RTSs is to accommodate testing of the UARS propulsion module (PM) in
preparation for the UARS orbit adjust maneuver later this year.
Testing the command side of the I/P Net remains on hold. Testing will
begin when MSOCC provides a command loop interface from the POCC through
the I/P Net to the UARS Simulator.
OTHER SIGNIFICANT EVENTS
Continued testing of the upgrade to the OBC Power Monitor (PMON) flight
software. PMON software testing continued this week. Multiple simulations
on the new AP software were conducted. However, in the course of testing,
several problems arose with the ground system software test packs. These
problems will be discussed with the programmer and hopefully corrected next
week.
Completed development and testing of the PMON update uplink procedure to
load the new PMON flight software. The procedure was successfully tested.
Work continues on the UARS orbit adjust planning re-scheduled for 12 August
1998. FOT was tasked to developed a command sheet and STOL procedure to
refill the B/D thruster lines. These were delivered to the customer on
Thursday afternoon following the 04 June 1998 meeting, and are pending
approval at the scheduled 09 June 1998 follow-up meeting.
The line refill procedure was tested with the simulator on 06 June 1998.
Minor adjustments to the UTTS were required to complete the simulation.
The procedure was successfully tested.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
A meeting to discuss the planned orbit adjust was held on 04 June 1998.
The following points of interest came out of the meeting:
The B/D attitude thrusters will be used, and hence the B/D fuel lines
will be refilled by an FOT-defined procedure. The procedure will be
reviewed by Code 570 at a meeting scheduled for 2:00 p.m. on 09 June
1998. The date for refilling the lines will be set at this meeting.
The B/D attitude and translation will be tested on 28 July 1998. The
thruster firings will be of very short duration, and will be analyzed
post-pass by FOT and FDF to determine if performance of the PM is nominal
for the fairly long-duration orbit adjust to follow.
The orbit adjust is now scheduled for 12 August 1998.
The representatives for the Lockheed Martin Spacecraft Control System - 21
(SCS-21) were at GSFC from 01 June through 05 June 1998 to discuss the UARS
functional requirements for a ground support system. The SCS-21 system is
being considered as a COTS upgrade system for the UARS command and
telemetry system. Meeting were conducted all week to discuss the
requirements in detail and to assess the SCS-21 current generic
capabilities and any special implementations required for UARS support.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 15 June 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
Appendix F, UARS Instrument Operations Summary
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 June 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
36709 151/0135 98-185 HRDI HRDI ON: HI Power
36732 152/1531 98-186 HRDI Microload
36737 152/2330 98-187 MLS Power OFF
36738 153/0107 98-188 HALOE Power ON
36765 154/2036 98-189 SUSIM Reloaded Microload
36771 155/0528 98-190 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 June 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 36788-97
NBBTAPEPOS R-LO NBTR B TAPE POSITION 36788
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 June 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
36771 155/05:33 (04 Jun 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 June 1998
30 May 1998 - GMT Day 150 Beta = 67.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -11.2 +3.87/+2.60 +21.2/-15.2 26.7 1.034 18.0
3 +84.0/ -5.6 +2.60/+1.66 +20.4/-15.2 26.7 1.031 17.5
05 Jun 1998 - GMT Day 156 Beta = 43.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.62 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -22.4 +5.47/+4.51 +28.4/-16.0 25.4 1.024 21.6
3 +84.0/ -11.2 +4.51/+3.23 +27.2/-15.2 25.4 1.028 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 June 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-027 FOT 36698 150/0819 N/A TAC #4 Crashed
98-028 FOT 36746 153/1307 14170 Generic Late Acquisition
98-029 SUSIM 36766 154/2036 N/A Incorrect Microload
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 Generic Late Acq for May 1998 (Open until 31 May 1998).
98-026 FOT 36680 149/0312 n/a TAC Crash
98-027 FOT 36698 150/0819 n/a TAC Crash
98-028 Generic Late Acq for June 1998 (Open until 30 June 1998).
98-029 SUSIM 36765 154/2036 n/a Incorrect micro load
_______________________________________________________________
APPENDIX F
UARS INSTRUMENT OPERATIONS SUMMARY
Week ending 05 June 1998
INSTRUMENT SUMMARY FOR MONTH OF May 1998 (hours)
UNPLANNED UNSCHEDULED
ON OFF OFF OFF
(Insufficient (Other than
Instrument (Actual) (Scheduled) Power) Power Problem) COMMENTS
---------- ------ ----------- ----------- ------------- ------------
ACRIM 744 0 0 0
HALOE 254 490 0 0
HRDI 022 722 0 0
MLS 335 409 0 0
PEM (axis) 197 338 209 0 Note 1
PEM (zeps) 116 628 0 0
PEM (neps) 0 744 0 0
SOLSTICE 744 0 0 0
SUSIM 744 0 0 0
WINDII 295 264 185 0 Note 1
Note 1 - Autonomous turn-off due to power balance
Cold side of spacecraft was facing northward: 4.5 days
Cold side of spacecraft was facing southward: 26.5 days
SEVNTFW23